OpenStudy (anonymous):

Find the inverse of the power function. f(x)=-10x^6 x>=0 What I got was... sixth root of x/-10

OpenStudy (nincompoop):

the inverse of a power is root

OpenStudy (anonymous):

huh?

OpenStudy (nincompoop):

the inverse of a power is root

OpenStudy (nincompoop):

x^2 = x^(1/2) or sqrt(x)

OpenStudy (nincompoop):

f(x)=-10x^6 y = -10x^6 solve for x

OpenStudy (anonymous):

im still lost

OpenStudy (nincompoop):

do you know how to solve for the x?

OpenStudy (nincompoop):

y = -10 * x^6 can you do it now?

OpenStudy (anonymous):

yea i know how to solve for x but my book says to solve for y when you are doing inverse

OpenStudy (nincompoop):

solve for the x first then switch the x and y at the end

OpenStudy (anonymous):

it'd be the same thing that i got.

OpenStudy (nincompoop):

then you are correct

OpenStudy (nincompoop):

have more confidence in your ability solve problems instead of looking for people's approval

OpenStudy (anonymous):

then where does the x>=0 come in?

OpenStudy (nincompoop):

domain

OpenStudy (anonymous):

is it kinda supposed to be like  and then i need to use x>=0 to fnd out whether its positive or negative?